OpenStudy (anonymous):

find the derivative x^3-x^2/28+8x+3

OpenStudy (anonymous):

\[\frac{x^3 - x^2}{28 + 8x + 3}\] ??

OpenStudy (anonymous):

thats the final answer?

OpenStudy (anonymous):

No, I'm asking if that is what you mean

OpenStudy (anonymous):

mine is correct

OpenStudy (anonymous):

Please do not simply post wolfram links as answers only. It's important to verify that the OP has indeed written the equation correctly (with parenthesis and such in the appropriate places)

OpenStudy (anonymous):

Secondly, without any meaningful dialogue, wolfram is not likely to help the OP.

OpenStudy (anonymous):

\[x ^{3}-(x ^{2}/28)+8x+3\]

OpenStudy (anonymous):

What, in particular is stumping you. (Here, x^2 / 28 is the same as (1/28)x)

OpenStudy (anonymous):

sorry: (1/28)x^2

OpenStudy (lgbasallote):

the question is @qpHalcy0n is it "help" the "OP" is really asking for?

OpenStudy (anonymous):

Which is why I'm not going to give any free hand-outs ;]

OpenStudy (anonymous):

i just have been working on this problem for awhile and can't seem to come up with the right answer and i dont know why

OpenStudy (anonymous):

Ok, what ARE you coming up with, then?

OpenStudy (anonymous):

y=3x^2-(x/14)+8

OpenStudy (anonymous):

Well, that's the correct solution. So, where's the problem?

OpenStudy (anonymous):

good question...its for an online assignment and wont be accepted as correct

OpenStudy (anonymous):

I detest online math assignments. Presuming you wrote it entirely correctly, then that is the correct solution. It MAY want you to say (1/14)x for that term.

OpenStudy (anonymous):

so then how should it look?

OpenStudy (anonymous):

The way you posted it is completely valid: \[3x^2 - \frac{1}{14}x + 8\] or \[3x^2 - \frac{x}{14} + 8\] Are both valid.

OpenStudy (anonymous):

the first one is correct thanks

OpenStudy (anonymous):

Wow, that's retarded. Exactly why I hate online math assignments ;]

OpenStudy (anonymous):

exactly and thats no the first time its happened either

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

I believe it. Tricking good work into thinking you're wrong.

OpenStudy (anonymous):

\[\frac{ d }{ dx } \left( \frac{ x^3 - x^2 }{ 28+8x+3 }\right)\] you mean this one ?

OpenStudy (anonymous):

No. We got it.